# Measurement of the longitudinal flow-plane decorrelation using multi-plane cumulants in $\sqrt{s_{{}_{\mathrm{NN}}}}$ = 200 GeV Au+Au, Ru+Ru, and Zr+Zr collisions

### v1

@@ II.1 Datasets and Event Selection
Figure 1: Diagram illustrating four rapidity regions used for $T_{n}$ calculations.
The primary dataset used in this work was collected by the STAR experiment et al. (1994) for isobar Ru+Ru and Zr+Zr collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV in $2018$ to search for the Chiral Magnetic Effect (CME) et al. (2022). The main detectors used in the analysis are the Time Projection Chamber (TPC) et al. (STAR Collaboration), Event Plane Detector (EPD) et al. (STAR Collaboration), and Beam-Beam Counter (BBC) et al. (STAR Collaboration). To study the system size dependence of the $T_{n}$ observable, we also analyze the data of Au$+$Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV taken in years $2011$ and $2016$. The EPD subsystem used in the analysis of the isobar collisions is an detector upgrade installed in the STAR detector system in $2018$. It significantly improves the measurements of the event-plane angle in forward and backward pseudorapidity regions, compared to the BBC detector used earlier. Since the Au+Au collisions data in $2011$ and $2016$ were taken using the TPC and BBC, the BBC is used in the analyses of these datasets to estimate the event-plane orientations at forward and backward pseudorapidities. The results using Au+Au and isobar collisions are then corrected by corresponding BBC or EPD event plane resolutions.
@@
Figure 2: Left: The second harmonic event plane resolutions from the TPC, EPD, and BBC detectors. Right: The third harmonic sub-event plane resolutions from the TPC and EPD. The filled and open data points are for Ru$+$Ru and Zr$+$Zr collisions, respectively. The Zr$+$Zr data points are shifted towards right along the x-axis for clarity.
Many details of this analysis are similar to those in the CME isobar analysis. The event and the track selections, as described in details in Ref. et al. (2022), have been used on the same datasets for this analysis. The vertex distribution on isobar collisions exhibited asymmetry with a peak around $5$ cm due to online vertex selection. Therefore, an asymmetric vertex position cut of $-35<v_{z}<25$ cm on the primary $z-$vertex measured using a coordinate system with the origin at the TPC center was applied to maximize the statistics. Additional cut of $v_{r}=\sqrt{v_{x}^{2}+v_{y}^{2}}<2$ cm in the transverse plane was also applied for each collision event. In order to ensure the selection of good events, requirement of $|v_{z,\mathrm{TPC}}-v_{z,\mathrm{VPD}}|<5$ cm was imposed on the primary collision vertices, where $v_{z,\mathrm{VPD}}$ represents the primary interaction position along the beam direction measured by the Vertex Position Detectors (VPDs) et al. (2004). In addition, we require tracks reconstructed in the TPC to have a distance of closest approach (DCA) to the primary vertex to be less than $3$ cm. We also require each track to have greater than $15$ ionization points reconstructed in the TPC.
The symmetric vertex position cuts of $-30<v_{z}<30$ cm and $-6<v_{z}<6$ cm were applied to the $2011$ and $2016$ Au+Au collision datasets, respectively, as online vertex reconstruction for these collisions indicated symmetric $v_{z}$ distributions. Due to the deployment of the Heavy Flavor Tracker (HFT) Bouchet (2009) as an upgrade to the inner tracking system of the STAR experiment, a relatively narrow $v_{z}$ cut was applied for the $2016$ Au+Au collision dataset. Other track selection cuts used for the Au+Au collision analysis are same as those used for the isobar collision analysis.
@@ II.2 Event Planes Reconstruction
The second- and third-order flow vectors $Q_{n}\equiv(Q_{n,x},Q_{n,y})$ and the event-plane angles $\Psi_{n}$ are reconstructed from the anisotropic flow itself following the procedure described in Refs. Poskanzer and Voloshin (1998); Voloshin et al. (2010):
@@
where $w_{i}$ is the weight for $i^{th}$ particle, and the sum goes over all particles used in the calculation. For the flow vector and event-plane angle reconstruction in the TPC, primary tracks with $2.00.2<p_{\mathrm{T}}<$ 2.0 GeV/$c$ were taken from $-1<\eta<-0.1$ and $10.1<\eta<$ 1. The flow vectors were reconstructed with unit weight for the Particle-level calculations and with transverse momentum weight for the Q-level calculations (see below). Charged particles emitted in the forward and backward rapidity regions create a signal distribution with distinct peaks, each representing different numbers of minimally ionizing particles (nMIP) in the EPD, BBC tiles. In the EPD, the event-plane angles were reconstructed using nMIP of each EPD tile $i$ which satisfies $0.3\leq$ nMIP $\leq 3$ as weight, $w_{i}$, in Eqs. 7 and  8. The weight is set to $3$ if nMIP $>3$. Similar information is used to reconstruct event-plane angles in the BBC.
@@
Figure 2 shows the event-plane resolutions as a function of centrality for the second- and third-order anisotropic flow calculated for the TPC, EPD, and BBC pseudorapidity subevents in the isobar collisions. The event-plane resolution calculations were done with the two- and three-subevent methods Voloshin et al. (2010).
II.3 Calculation of the Four-Plane Correlator
@@
Figure 3: (Color online) Resolution ${\rm Res}(T_{n}\{ba;dc\})$ plots for the Q-level (left) and Particle-level (right) calculations using Ru$+$Ru (solid markers) and Zr$+$Zr (open markers) collisions for the second-order anisotropic flow. The Zr$+$Zr data points are shifted towards right along x-axis for clarity.
In this work, the measurements of the $T_{n}\{ba;dc\}$ correlators for the second- and third-order anisotropic flow are performed using flow-plane angles reconstructed in the TPC and EPD (or BBC) subsytems of the STAR detector. Therefore, in the measurements presented in this paper, the pseudorapidity windows $a$, $b$, $c$, and $d$ from the backward, mid-, and forward pseudorapidity regions as depicted in Fig. 1 correspond to the EPD-east ($-5.1<\eta<-2.1$) or BBC-east ($-5.0<\eta<-3.3$), TPC-east ($-1.0<\eta<-0.1$), TPC-west ($1.00.1<\eta<$ 1.0), and EPD-west ($5.12.1<\eta<$ 5.1) or BBC-west ($5.03.3<\eta<$ 5.0) regions, respectively.
@@
The measurement using four flow planes from the east and west regions of the TPC and EPD (or BBC) is referred as the Q-level calculation, which follows Eq. 10. Similarly, the measurement using two flow planes from the east and west regions of the EPD (or BBC) along with the particle track’s azimuth $\phi$ from the east and west regions of the TPC is identified as the Particle-level calculation. The Particle-level calculation follows Eq. 11 for the $T_{n}\{ba;dc\}$ observable.
Experimentally, the four-plane correlator observable, $T_{n}$, can be calculated as
[…]
for the Q-level calculation and
[…]
for the Particle-level calculation.
The observed cumulant $T^{obs}_{n}\{ba;dc\}$ in the numerator of Eq. 10 is calculated using four $n^{th}$-harmonic event planes from the backward, mid-, and forward pseudorapidity regions $a$, $b$, $c$, and $d$ as shown in Fig. 1, which has to be corrected for the event-plane resolution. The term ${\rm Res}(T_{n}\{ba;dc\})$ in the denominator is used to correct $T^{obs}_{n}\{ba;dc\}$ measurement for the event-plane resolution.
@@
Figure 5: (Color online) Centrality dependence of different terms contributing to $T^{obs}_{2}\{ba;dc\}$, as given in Eqs. 19 and 20: Solid markers show Ru$+$Ru and open markers Zr$+$Zr results for the Q-level (left panel) and Particle-level (right panel) calculations. The Zr$+$Zr data points are shifted towards right for clarity.
@@
Since Eq. 10 utilizes the event planes reconstructed from the four different pseudorapidity regions, ${\rm Res}(T_{n}\{ba;dc\})$ in general can be expressed as the product of the four event-plane resolutions,
@@
In practice, the analysis estimates the ${\rm Res}(T_{n}\{ba;dc\})$ term using formula
@@
for the Q-level calculations by identifying the following correlations
@@
as the product of the two event-plane resolutions. Using Eqs. 14 and 15, the product of the four event-plane resolutions in Eq. 12 follows Eq. 13. In the Particle-level calculations, the event-plane resolution correction ${\rm Res}(T^{p-level}_{n}\{ba;dc\})$ is estimated using azimuthal angles ($\phi^{b}$, $\phi^{c}$) of tracks in mid-pseudorapidity windows $b$ and $c$ instead of $\Psi^{b}_{n}$ and $\Psi^{c}_{n}$ in Eq. 13, i.e., using formula
@@
The set of plots in Fig. 3 shows the centrality dependence of ${\rm Res}(T_{n}\{ba;dc\})$ and ${\rm Res}(T^{p-level}_{n}\{ba;dc\})$ for the elliptic flow in isobar collisions. A similar set of plots for triangular flow in isobar collisions are shown in Fig. 4.
In the following, the centrality dependence of the observed four-plane cumulant $T^{obs}_{n}\{ba;dc\}$ and its contributing components are discussed. The $T^{obs}_{n}\{ba;dc\}$ correlator, explicitly derived in Ref. Xu et al. (2022), consists of three terms:
$\begin{split}&T^{obs}_{n}\{ba;dc\}\\ &=\langle\langle\sin[n(\Psi^{b}_{n}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\Psi^{c}_{n})]\rangle\rangle\\ &\equiv\langle\sin[n(\Psi^{b}_{n}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\Psi^{c}_{n})]\rangle\\ &-\frac{1}{2}\langle\cos[n(\Psi^{b}_{n}-\Psi^{d}_{n})]\rangle\langle\cos[n(\Psi^{a}_{n}-\Psi^{c}_{n})]\rangle\\ &+\frac{1}{2}\langle\cos[n(\Psi^{a}_{n}-\Psi^{d}_{n})]\rangle\langle\cos[n(\Psi^{b}_{n}-\Psi^{c}_{n})]\rangle,\end{split}$ (19)
[…]
$\begin{split}&(T^{obs}_{n}\{ba;dc\})^{p-level}\\ &=\langle\langle\sin[n(\phi^{b}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\phi^{c})]\rangle\rangle\\ &\equiv\langle\sin[n(\phi^{b}-\Psi^{a}_{n})]\sin[n(\Psi^{d}_{n}-\phi^{c})]\rangle\\ &-\frac{1}{2}\langle\cos[n(\phi^{b}-\Psi^{d}_{n})]\rangle\langle\cos[n(\Psi^{a}_{n}-\phi^{c})]\rangle\\ &+\frac{1}{2}\langle\cos[n(\Psi^{a}_{n}-\Psi^{d}_{n})]\rangle\langle\cos[n(\phi^{b}-\phi^{c})]\rangle,\end{split}$ (20)
Figures 5 and 6 show the centrality dependence of different terms contributing to the $T^{obs}_{n}\{ba;dc\}$ observable for isobar collisions. Results from the Q-level and Particle-level calculations (as discussed in Sec. II.3) are presented side-by-side for better comparison. The centrality dependence of different terms from the Q-level and Particle-level approaches is slightly different. This effect is similar to the one responsible for the difference in the centrality dependence of the event-plane resolution and elliptic flow.
@@
Figure 7: Ratio observable $r_{n}$ (defined in Eq. 22) as a function of centrality for the second- (left) and third-order (right) anisotropic flow. (Color online) Red circle and blue star data points correspond, respectively, to Ru$+$Ru and Zr$+$Zr collisions. The systematic uncertainties are indicated by boxes. The Zr+Zr data points are shifted towards right for clarity.

### v2

@@ II.1 Datasets and Event Selection
The primary dataset used in this work was collected by the STAR experiment et al. (1994) for isobar Ru+Ru and Zr+Zr collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV in $2018$ to search for the Chiral Magnetic Effect (CME) et al. (2022). The main detectors used in the analysis are the Time Projection Chamber (TPC) Collaboration) (2003), Event Plane Detector (EPD) Collaboration) (2020), and Beam-Beam Counter (BBC) Collaboration) (2008). To study the system size dependence of the $T_{n}$ observable, we also analyze the data of Au++Au collisions at $\sqrt{s_{{}_{\mathrm{NN}}}}=200$ GeV taken in years $2011$ and $2016$. The EPD subsystem used in the analysis of the isobar collisions is an detector upgrade installed in the STAR detector system in $2018$. It significantly improves the measurements of the event-plane angle in forward and backward pseudorapidity regions, compared to the BBC detector used earlier (see Fig. 2). Since the Au+Au collisions data in $2011$ and $2016$ were taken using the TPC and BBC, the BBC is used in the analyses of these datasets to estimate the event-plane orientations at forward and backward pseudorapidities. The results using Au+Au and isobar collisions are then corrected by corresponding BBC or EPD event-plane resolutions.
@@
Figure 2: Left: The second harmonic event plane resolutions from the TPC, EPD, and BBC detectors. Right: The third harmonic sub-event plane resolutions from the TPC and EPD. The filled and open data points are for Ru++Ru and Zr++Zr collisions, respectively. The Zr++Zr data points are shifted towards right along the x-axis for clarity. The results shown are from the two-subevent method.
Many details of this analysis are similar to those in the CME isobar analysis. The event and the track selections, as described in details in Ref. et al. (2022), have been used on the same datasets for this analysis. The vertex distribution on isobar collisions exhibited asymmetry with a peak around $5$ cm due to online vertex selection. Therefore, an asymmetric vertex position cut of $35<v_{z}<25$ cm on the primary $z$-vertex measured using a coordinate system with the origin at the TPC center was applied to maximize the statistics. Additional cut of $v_{r}=\sqrt{v_{x}^{2}+v_{y}^{2}}<2$ cm in the transverse plane was also applied for each collision event. In order to ensure the selection of good events, requirement of $|v_{z,\mathrm{TPC}}-v_{z,\mathrm{VPD}}|<5$ cm was imposed on the primary collision vertices, where $v_{z,\mathrm{VPD}}$ represents the primary interaction position along the beam direction measured by the Vertex Position Detectors (VPDs) et al. (2004). In addition, we require tracks reconstructed in the TPC to have a distance of closest approach (DCA) to the primary vertex to be less than $3$ cm. We also require each track to have greater than $15$ ionization points reconstructed in the TPC.
The symmetric vertex position cuts of $30<v_{z}<30$ cm and $6<v_{z}<6$ cm were applied to the $2011$ and $2016$ Au+Au collision datasets, respectively, as online vertex reconstruction for these collisions indicated symmetric $v_{z}$ distributions. Due to the deployment of the Heavy Flavor Tracker (HFT) Bouchet (2009) as an upgrade to the inner tracking system of the STAR experiment, a relatively narrow $v_{z}$ cut was applied for the $2016$ Au+Au collision dataset. Other track selection cuts used for the Au+Au collision analysis are same as those used for the isobar collision analysis.
@@ II.2 Event Planes Reconstruction
The second- and third-order flow vectors $Q_{n}\equiv(Q_{n,x},Q_{n,y})$ and the event-plane angles $\psi_{n}$ (an estimate of the flow-plane angle $\Psi_{n}$) are reconstructed from the anisotropic flow itself following the procedure described in Refs. Poskanzer and Voloshin (1998); Voloshin et al. (2010):
@@
where $w_{i}$ is the weight for $i^{th}$ particle, and the sum goes over all particles used in the calculation. For the flow vector and event-plane angle reconstruction in the TPC, primary tracks with $0.2<p_{\mathrm{T}}<2.0$ GeV/$c$ were taken from $1<\eta<-0.1$ and $0.1<\eta<1$. The flow vectors were reconstructed with unit weight for the Particle-level calculations and with transverse momentum weight for the Q-level calculations (see below). Charged particles emitted in the forward and backward pseudorapidity regions create a signal distribution with distinct peaks, each representing different numbers of minimally ionizing particles (nMIP) in the EPD and BBC tiles. In the EPD, the event-plane angles were reconstructed using nMIP of each EPD tile $i$ which satisfies $0.3\leq$ nMIP $\leq 3$ as weight, $w_{i}$, in Eqs. 8 and  9. The weight is set to $3$ if nMIP >3>3 to suppress distortions from large Landau fluctuations in the energy-loss distribution. Similar information is used to reconstruct event-plane angles in the BBC.
@@
Figure 2 shows the event-plane resolutions ${\rm Res}(\Psi_{n})\equiv\langle\cos[n(\psi_{n}-\Psi_{n})]\rangle$ as a function of centrality for the second- and third-harmonic anisotropic flow calculated for the TPC, EPD, and BBC pseudorapidity subevents in the isobar collisions. The event-plane resolution calculations were done with the two- and three-subevent methods Voloshin et al. (2010). Figure 2 presents calculations using the two-subevent method. Note that the calculation of the event-plane resolutions using rapidity subevents might be affected by flow-plane decorrelations. See the next section for more details.
@@
Figure 3: (Color online) Resolution ${\rm Res}(T_{n}\{ba;dc\})$ plots for the Q-level (left) and Particle-level (right) calculations in Ru++Ru (solid markers) and Zr++Zr (open markers) collisions for the second-order anisotropic flow. The Zr++Zr data points are slightly shifted along x-axis for clarity.
@@
II.3 Calculation of the Four-Plane Correlator
In this analysis, the measurements of the $T_{n}\{ba;dc\}$ correlators for the second- and third-order anisotropic flow are performed using event-plane angles, $\psi_{n}$’s, reconstructed in the STAR TPC and EPD (or BBC) detectors. The pseudorapidity windows $a$, $b$, $c$, and $d$, depicted in Fig. 1, correspond to the EPD-east ($5.1<\eta<-2.1$) or BBC-east ($5.0<\eta<-3.3$), TPC-east ($1.0<\eta<-0.1$), TPC-west ($0.1<\eta<1.0$), and EPD-west ($2.1<\eta<5.1$) or BBC-west ($3.3<\eta<5.0$) regions, respectively.
The measurement using four event planes, from the east and west regions of the TPC and EPD (or BBC), is referred to as the Q-level calculation, and the measurement using two event planes from the east and west regions of the EPD (or BBC) and the particle’s azimuth $\phi$ from the east/west regions of the TPC is identified as the Particle-level calculation.
@@ II.3 Calculation of the Four-Plane Correlator
Figure 5: (Color online) Centrality dependence of different terms contributing to $T^{obs}_{2}\{ba;dc\}$, as given in Eqs. 12 and 18: Solid markers show Ru++Ru and open markers Zr++Zr results for the Q-level (left) and Particle-level (right) approaches. The Zr++Zr data points are shifted towards right for clarity.
@@
Experimentally, the four-plane correlator $T_{n}$ is calculated as
$T_{n}\{ba;dc\}(\mathrm{Q\,\text{-}\,level})=\frac{T_{n}^{obs}\{ba;dc\}}{\rm Res{\it(T_{n}\{ba;dc\})}},$ (11)
where the term ${\rm Res}(T_{n}\{ba;dc\})$ in the denominator is used to correct $T^{obs}_{n}\{ba;dc\}$ for the event-plane resolution, and discussed below. The numerator, the “observed” $T_{n}^{obs}$ correlator, for Q-level calculations is evaluated following the $T_{n}$ definition, Eq. 5:
$\begin{split}&T^{obs}_{n}\{ba;dc\}\\ &\equiv\langle\langle\sin[n(\psi^{b}_{n}-\psi^{a}_{n})]\sin[n(\psi^{d}_{n}-\psi^{c}_{n})]\rangle\rangle\\ &=\langle\sin[n(\psi^{b}_{n}-\psi^{a}_{n})]\sin[n(\psi^{d}_{n}-\psi^{c}_{n})]\rangle\\ &-\frac{1}{2}\langle\cos[n(\psi^{b}_{n}-\psi^{d}_{n})]\rangle\langle\cos[n(\psi^{a}_{n}-\psi^{c}_{n})]\rangle\\ &+\frac{1}{2}\langle\cos[n(\psi^{a}_{n}-\psi^{d}_{n})]\rangle\langle\cos[n(\psi^{b}_{n}-\psi^{c}_{n})]\rangle.\end{split}$ (12)
Where, $\psi^{i}_{n}\textrm{ (for }i=a,b,c,d)$ are the event-plane angles of $n^{\rm th}$-order harmonic flow. All the three terms contributing to the $T^{obs}_{n}\{ba;dc\}$ observable are calculated independently in this analysis for the Q-level and Particle-level methods.
Since $T_{n}$ measurements utilizes the event planes reconstructed from four different pseudorapidity regions, ${\rm Res}(T_{n}\{ba;dc\})$ can be expressed as the product of the four event-plane resolutions,
@@
In practice, we estimate ${\rm Res}(T_{n}\{ba;dc\})$ using formula
@@
and identifying the following correlations
@@
as the product of the two event-plane resolutions. Using Eqs. 15 and 16, the product of the four event-plane resolutions in Eq. 13 follows Eq. 14. The calculations of the event-plane resolution from correlation of two event planes (or three event planes in three sub-event method) is affected by the flow-plane decorrelations. This effect is relatively small, of the order of 5–15% based on the previous and our measurements of the ratio correlator (see e.g. Fig. 7 below), and significantly smaller than other uncertainties in the measurements. Note that accounting for this effect would slightly decrease the values of $T_{n}$ reported below.
In the Particle-level calculations, the $T_{n}^{obs}$ correlator and the event-plane resolution correction ${\rm Res}(T^{p-level}_{n}\{ba;dc\})$ are calculated using azimuthal angles ($\phi^{b}$, $\phi^{c}$) of tracks in mid-pseudorapidity windows $b$ and $c$ instead of $\psi^{b}_{n}$ and $\psi^{c}_{n}$, i.e., using formulae
$\begin{split}T_{n}\{ba;dc&\}(\mathrm{Particle\,\text{-}\,level})\\ &=\frac{(T_{n}^{obs}\{ba;dc\})^{p-level}}{\rm Res{\it(T^{p-level}_{n}\{ba;dc\})}},\end{split}$ (17)
$\begin{split}&(T^{obs}_{n}\{ba;dc\})^{p-level}\\ &\equiv\langle\langle\sin[n(\phi^{b}-\psi^{a}_{n})]\sin[n(\psi^{d}_{n}-\phi^{c})]\rangle\rangle\\ &=\langle\sin[n(\phi^{b}-\psi^{a}_{n})]\sin[n(\psi^{d}_{n}-\phi^{c})]\rangle\\ &-\frac{1}{2}\langle\cos[n(\phi^{b}-\psi^{d}_{n})]\rangle\langle\cos[n(\psi^{a}_{n}-\phi^{c})]\rangle\\ &+\frac{1}{2}\langle\cos[n(\psi^{a}_{n}-\psi^{d}_{n})]\rangle\langle\cos[n(\phi^{b}-\phi^{c})]\rangle,\end{split}$ (18)
and
@@
Figures. 3 and 4 show the centrality dependence of ${\rm Res}(T_{n}\{ba;dc\})$ and ${\rm Res}(T^{p-level}_{n}\{ba;dc\})$ for the elliptic and triangular flow in isobar collisions.
The $T^{obs}_{n}\{ba;dc\}$ correlator consists of three terms, and the standard error propagation method is used to calculate the statistical uncertainties on the final observable. The uncertainties on the second and third terms in Eqs. 12 and 18 are negligible compared to that on the first term (see Figs. 5 and 6). Hence, the uncertainties on the final observable comes predominantly from the first term even though the three terms in Eqs. 12 and 18 are correlated.
Figures 5 and 6 show the centrality dependence of different terms contributing to $T^{obs}_{n}\{ba;dc\}$ for isobar collisions. Results from the Q-level and Particle-level calculations are presented side-by-side for better comparison. The centrality dependence of different terms in two approaches are slightly different. This difference is similar to the one responsible for the difference in the centrality dependence of the event-plane resolution and elliptic flow.
@@
Figure 7: (Color online) Ratio observable $r_{n}$ (Eq. 23) as a function of centrality for the second- (left) and third-harmonic (right) anisotropic flow. Red circles and blue stars correspond to Ru++Ru and Zr++Zr collisions, respectively. The systematic uncertainties are indicated by boxes. The Zr+Zr data points are shifted towards right for clarity.